OpenStudy (einstein):

The lifetime of a device behaves according to the probability law P(t, infinity) = 1/(1+t) , t >0 a. calculate the probability that the lifetime of the device is greater or equal than 3 but less than 8 b. Calculate the probability that the device last at least another 5 hours given that it has already lasted 5.

OpenStudy (anonymous):

ok that was way way wrong, sorry

OpenStudy (anonymous):

embarrassingly wrong

OpenStudy (anonymous):

but fortunately the answer is much much easier the probability that the life is greater than 3 is \(p(3)=\frac{1}{1+3}=\frac{1}{4}\) the probability it is greater than 8 is \(p(8)=\frac{1}{9}\) so the probability it is greater than 3 and less than 9 is \(\frac{1}{4}-\frac{1}{9}\)

OpenStudy (anonymous):

second one \[P(A|B)=\frac{P(A\cap B)}{P(B)}\] \[P(B)=p(5)=\frac{1}{6}\] and \[P(A \cap B)=\frac{1}{6}-\frac{1}{11}\]

OpenStudy (einstein):

Here probability is a continuous function of time and you are applying discrete formulas

OpenStudy (anonymous):

formula is the same

OpenStudy (einstein):

oh ya, thaks ,good work

OpenStudy (einstein):

I also have a probability problem would you like to solve that Mr. satellite73